(1/3)y to the second power + 12= 5x. Where. X=3.
Answer by ewatrrr(24785)


Hi,
(1/3)y^2 + 12 = 5x
substituting x = 3
(1/3)y^2 + 12 = 15
Multiplying each side of the equation by 3
y^2 +36 = 45
y^2 = 9
y = 3
checking our answer
9/3 + 12 = 3 + 12 = 15 = 5*3